OpenStudy (anonymous):

solve find 10% of 35, find 25% of 80

OpenStudy (anonymous):

1)10%0f 35=35*(10/100)=35/10=3.5 2)25%of80=80*(25/100)=80/4=20

OpenStudy (anonymous):

when u wrote the problem 35w what is the w for ?

OpenStudy (anonymous):

multiply ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how dose 35 times 10/100 = to the answer u got ex-plane please if u don't mind.

OpenStudy (anonymous):

1 percentage of a number means, 1 in 100th of the given number(1/100) 10 percentage of 35,means ,10 in 100th of 35 ie 35 x 10/100

OpenStudy (anonymous):

thanks